Rh=1/resistivity

Why hall coefficient is equal to 1/restivity why this relation come from

what is the relation between resistivity and hall coefficient may I get the relation with proof



1
Expert's answer
2021-02-04T10:20:07-0500

In wires, electrons instead of holes are flowing, so "vx → −vx \space and \space q → −q. \space \space Also \space \space Ey = −
VH
/
w"

"{\displaystyle V_{\mathrm {H} }=v_{x}B_{z}w}"

The conventional "hole" current is in the negative direction of the electron current and the negative of the electrical charge which gives "Ix = ntw(\u2212vx)(\u2212e)"  where n is the charge carrier density, tw is the cross-sectional area, and −e is the charge of each electron. Solving for "{\displaystyle w}" and plugging into the above gives the Hall voltage:

"{\displaystyle V_{\mathrm {H} }={\frac {I_{x}B_{z}}{nte}}}"

If the charge build-up had been positive (as it appears in some semiconductors), then the VH assigned in the image would have been negative (the positive charge would have built up on the left side).

The Hall coefficient is defined as

"{\displaystyle R_{\mathrm {H} }={\frac {E_{y}}{j_{x}B_{z}}}}
\space or \space 
{\displaystyle E=-R_{H}(J_{c}\times B)}{\displaystyle E=-R_{H}(J_{c}\times B)}"

where j is the current density of the carrier electrons, and Ey is the induced electric field. In SI units, this becomes

"{\displaystyle R_{\mathrm {H} }={\frac {E_{y}}{j_{x}B}}={\frac {V_{\mathrm {H} }t}{IB}}={\frac {1}{ne}}.}"

(The units of RH are usually expressed as m3/C, or Ω·cm/G, or other variants.) As a result, the Hall effect is very useful as a means to measure either the carrier density or the magnetic field.